OpenStudy (anonymous):

how would you do 3X+5y=0 x-4y=0

OpenStudy (anonymous):

Is this a system of equations?

OpenStudy (anonymous):

its substitution method

OpenStudy (anonymous):

x-4y=0 +4y +4y x=4y 3(4y)+5y=0 12y+5y=0 17y=0 divide each side by 17 y=0

OpenStudy (anonymous):

x=4y x=4(0) x=0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

could you help me with more questions??

OpenStudy (anonymous):

4k+m=9 3k-4m=2

OpenStudy (anonymous):

yep one sec

OpenStudy (anonymous):

4k+m=9 -4k -4k m=-4k+9 3k-4(-4k+9)=2 3k+16k-36=2 +36 +36 19k=38 divide each side by 19 k=2

OpenStudy (anonymous):

4(2)+m=9 8+m=9 -8 -8 m=1

OpenStudy (anonymous):

Do you understand what I did to get those answers or would you like me to explain?